OpenStudy (anonymous):

simplify 45 sqrt * 80 sqrt please help me solve with working

OpenStudy (mertsj):

\[\sqrt{45}\sqrt{80}=\sqrt{9(5)}\sqrt{16(5)}=3\sqrt{5}(4\sqrt{5})=12(5)=60\]

OpenStudy (anonymous):

Is this what you mean? \[\sqrt{45}*\sqrt{80}\]

OpenStudy (anonymous):

or \[\sqrt{45 *\sqrt{80}}\]